what is ther standard form of y=16-x+3x.
at the end of three in my equation i wanted to say 3 squared to the second power ,but it wouldnt let me type it so can you do it.

Answer by ewatrrr(24785)

y= 3x^2 - x + 16
y = 3(x-1/6)^2 - 1/12 + 16
(1/3)(y - 191/12) = (x- 1/6)^2